# Artificial Intelligence in Elementary STEM Education: A Systematic Review of Current Applications and Future Challenges

### v1

@@ 3.12.7 Privacy, Ethics, and Equity Considerations
Sensor-rich MR systems collecting granular behavioral data from elementary students raise substantial privacy concerns requiring proactive governance [37, 35]. Critical scholars warn that immersive technologies risk creating surveillance infrastructures if data collection, retention, and usage policies lack transparency and community oversight. Recommendations include limiting data collection to pedagogically essential streams, anonymizing identifiable information through pseudonymization and aggregation, obtaining informed parental consent with clear communication about data purposes, and designing analytic outputs as interpretable teaching tools rather than student monitoring systems [37, 35, 33].
@@ 3.12.8 Research Priorities and Knowledge Gaps
The field faces three critical knowledge gaps demanding targeted research investment. First, an empirical gap exists in validated XR implementations specifically designed for and evaluated with elementary STEM learners [28, 29]. Most rigorous effectiveness trials target adults or secondary students; evidence for elementary populations—particularly for integrated cross-disciplinary STEM learning—remains thin.
Second, a methodological gap persists in training datasets for machine learning models: publicly available multimodal behavioral datasets from children using XR systems for STEM learning do not exist, limiting researchers’ ability to develop and validate analytic algorithms or investigate generalizability across device platforms [34, 24].
Third, a pedagogical gap concerns adaptive scaffolding and teacher orchestration: research documenting effective real-time adaptation strategies based on behavioral analytics, and teacher-facing dashboard designs supporting instructional decision-making in XR-enhanced classrooms, remains urgently needed [31, 38].
@@ 3.12.9 Practical Recommendations for Pilot Implementations
Institutions considering XR pilots for elementary STEM should follow staged implementation guidance [28, 25]. Begin with small-scale deployments (1-2 classrooms, 4-6 week instructional units) aligned closely with existing STEM curricula, co-designing activities with classroom teachers to ensure pedagogical coherence. Implement clear behavioral signal tracking from initial pilots to enable model validation and iterative refinement [24]. Prioritize interpretable analytics—visualizations teachers can understand and act upon—while maintaining transparent data governance with documented consent, storage, and usage policies [35, 37]. Evaluate through convergent mixed methods combining quantitative metrics with teacher interviews and classroom observations, sharing findings through practitioner networks and research publications to build the community knowledge base [29, 28].
Critical gap: XR/spatial-computing technologies provide compelling theoretical affordances for embodied, spatially-grounded integrated STEM learning in elementary contexts, with modern devices generating rich behavioral data streams enabling sophisticated learning analytics. However, the evidence base remains dominated by health professions training and university implementations; rigorous empirical validation specifically targeting elementary STEM classrooms—particularly for cross-disciplinary integrated learning supported by conversational AI—is lacking. Moreover, critical privacy, equity, and pedagogical integration challenges require substantial investigation before XR can move from promising prototype to validated, equitably-deployed elementary STEM learning infrastructure [19, 20, 33, 25, 21, 37].
@@ 3.13 Educational Robots and Teachable Agents: Learning-by-Teaching Paradigms for Elementary STEM
Educational robots and teachable agents represent a distinct embodied AI paradigm where elementary students learn by teaching, explaining concepts to artificial tutees rather than receiving instruction from AI tutors. This pedagogical inversion—positioning the child as expert and the AI as learner—activates metacognitive processes, explanatory elaboration, and self-monitoring that align with cooperative learning mechanisms documented to enhance elementary learning outcomes [39]. However, while cooperative peer learning shows robust meta-analytic evidence in elementary mathematics and the theoretical parallels between peer teaching and AI-tutee interaction are compelling, direct experimental evaluations of educational robots or teachable agents deployed in elementary STEM classrooms remain limited, necessitating careful synthesis of adjacent literatures to guide future empirical work [40, 41, 42].
@@ 3.13.1 Theoretical Foundation: Cooperative Learning as Analogue for Tutee Interactions
Meta-analytic synthesis of elementary mathematics programs demonstrates that designs incorporating cooperative learning produce meaningful achievement gains, establishing peer explanation and collaborative problem-solving as core mechanisms linking social interaction to cognitive growth [39]. When students articulate their reasoning to peers, they externalize thought processes, identify gaps in understanding, refine explanations through iterative dialogue, and consolidate knowledge through teaching acts. These mechanisms provide direct theoretical justification for learning-by-teaching interventions with AI tutees: if explaining to human peers enhances learning, then explaining to artificial agents designed to elicit similar explanatory discourse should produce analogous cognitive benefits.
The pedagogical value of teaching others extends beyond simple rehearsal. Effective peer teaching requires monitoring the tutee’s understanding, adjusting explanations based on feedback, and scaffolding learning through strategic questioning—all metacognitive and regulatory processes that deepen the tutor’s own comprehension [39]. Educational robots and teachable agents can be designed to systematically elicit these processes through programmed responses signaling confusion, requesting clarification, or demonstrating partial understanding, thereby creating structured opportunities for elementary students to engage in explanatory and metacognitive verbalization within STEM learning contexts.
@@ 3.13.2 Implementation Realities: Teacher Preparation and Classroom Feasibility
Elementary classrooms present unique implementation constraints fundamentally shaping technology deployment feasibility. Teachers are typically subject generalists rather than STEM specialists, exhibiting variable content knowledge and pedagogical confidence across mathematics, science, engineering, and computational domains [41, 42]. Professional development supporting elementary STEM technology integration must therefore address both content knowledge and pedagogical practice, recognizing that teacher STEM identity, beliefs about technology, and confidence in orchestrating novel activities mediate implementation fidelity and sustainability.
Research on preservice and in-service elementary teacher preparation for integrated STEM instruction emphasizes co-design approaches where teachers participate in developing activities, scripting robot prompts, and establishing classroom routines before deployment [41, 42]. This participatory design philosophy treats teachers as essential partners whose practical wisdom about classroom management, developmental appropriateness, and curricular alignment ensures that robot-tutee interventions integrate smoothly into instructional sequences rather than disrupting established routines. Studies documenting effective elementary STEM teacher preparation highlight the importance of modeling exemplar implementations, providing scripted facilitation guides, and creating professional learning communities where teachers share strategies for managing technological and pedagogical challenges [41, 42].
@@ 3.13.3 Measurement Framework: Constructs, Instruments, and Developmental Considerations
Rigorous evaluation of learning-by-teaching robot interventions in elementary contexts requires multimodal assessment spanning cognitive learning outcomes, interaction processes, socio-emotional effects, and implementation fidelity [40, 43]. Cognitive outcomes should employ curriculum-aligned measures with established validity evidence for the targeted grade range, following psychometric best practices for evidence-based validity and reliability reporting [40]. Recent work validating elementary mathematics assessment instruments illustrates necessary procedures: item analysis, internal consistency documentation, construct validity through correlations with established measures, and examination of measurement invariance across grade bands to ensure scores are interpretable equivalently for K-2 versus grades 3-5 students [40, 43].
Interaction process measures capture the explanatory and metacognitive mechanisms hypothesized to mediate learning gains when students teach robots. Systematic sampling and coding of student-robot dialogue should prioritize: (1) concept explanations where students articulate domain principles, (2) strategy statements describing problem-solving approaches, (3) metacognitive verbalizations reflecting planning, monitoring, and evaluation, and (4) scaffolding behaviors such as questioning and prompting [39]. These codes align directly with cooperative learning literature documenting that student talk—particularly high-quality elaborations—predicts achievement gains, providing mechanistic justification for treating verbal interaction quality as a primary mediator variable [39].
Socio-emotional and behavioral outcomes require validated screening instruments appropriate for elementary populations. The Social Skills Improvement System Performance Screening Guide (SSIS PSG) demonstrates meta-analytic support as an elementary screening tool detecting behavioral and academic concerns, making it suitable for monitoring unintended effects of robot interventions on peer dynamics or classroom behavior [44]. Studies documenting structural validity and measurement invariance of socio-emotional instruments across grade clusters emphasize the importance of selecting scales whose psychometric properties support valid interpretation across the elementary range [43].
Implementation fidelity measurement captures teacher professional development participation, classroom enactment logs documenting session completion and adherence to protocols, and teacher self-reported content confidence before and after implementation [41, 42]. Because teacher knowledge and receptivity mediate technology adoption in elementary settings, measuring PD dose, teacher beliefs, and fidelity enables researchers to model these factors as moderators of student outcomes and to identify conditions supporting successful implementation [41, 42].
@@ 3.13.4 Data Collection Protocol and Analytic Strategies
A comprehensive data collection protocol for educational robot-tutee studies in elementary STEM integrates baseline and outcome cognitive testing, behavioral and socio-emotional screening, observational sampling of student-robot discourse, teacher implementation logs, and system telemetry from robot interactions [40, 39]. Cognitive assessments should employ standardized or researcher-developed instruments with demonstrated validity, reporting pre-post effect sizes, growth metrics, and psychometric properties to support meta-analytic synthesis [40, 39].
Discourse sampling requires audio-recording student-robot interactions and applying structured coding schemes with established inter-rater reliability. Cooperative learning evidence justifies prioritizing explanatory and metacognitive codes as intermediate outcomes theoretically linked to learning gains [39]. System logs from robots provide complementary objective data: timestamped interaction events, session durations, task sequences, and robot prompts delivered. Synchronizing telemetry with audio transcripts enables triangulation—linking coded verbal behaviors to system states and timestamped learning events to support process-outcome models [39].
Analytic approaches should employ multilevel modeling reflecting the nested structure of elementary data (students within classrooms within schools), because interventions typically deploy at classroom level and teacher practices explain substantial outcome variance [41, 42]. Hierarchical models partition variance appropriately, estimate classroom-level moderators (PD fidelity, teacher STEM identity), and support generalization to broader populations. Mediation analyses testing whether interaction processes (explanation quality, metacognitive verbalization frequency) mediate the robot intervention’s effect on learning outcomes provide mechanistic evidence aligning with cooperative learning theory [39].
Grade-band stratification and tests of measurement invariance address developmental heterogeneity across elementary years. Because working memory, executive function, and metacognitive capacity evolve substantially from kindergarten through grade 5, analyses should examine whether treatment effects, process mechanisms, or measurement properties differ across K-2 versus grades 3-5 clusters [43, 45]. Executive function screening or age/grade covariates help contextualize metacognitive verbalization patterns and test for differential responsiveness by developmental level.
@@ 3.13.5 Staged Research Pathway: Feasibility, Efficacy, Effectiveness
A methodologically rigorous research program for educational robots in elementary STEM should progress through staged designs balancing scientific control with ecological validity [41, 42, 39]. Feasibility studies establish that robots integrate into classroom routines without excessive disruption, that teachers can implement activities with reasonable PD support, that audio/video capture and telemetry synchronization function reliably, and that instruments yield interpretable data. Teacher-led pilots allow iterative refinement of robot scripts, facilitation guides, and data collection procedures before scaling to formal trials [41].
Efficacy trials employ cluster randomization or well-matched quasi-experimental designs assigning classrooms or schools to intervention versus comparison conditions, with comprehensive pre-post measurement, process sampling, and fidelity monitoring. Power calculations should account for classroom-level assignment and intraclass correlation typical of elementary settings [39, 40]. Cooperative learning meta-analyses provide benchmarks for expected effect sizes and variance structures informing power analysis.
Effectiveness studies examine scalability and sustainability under real-world conditions, explicitly measuring teacher adoption patterns, PD scalability, and contextual moderators such as school learning models (in-person, hybrid, remote instruction) that influence technology access and supervision [46]. Pragmatic effectiveness trials prioritize external validity and generalizability, documenting implementation variation and subgroup effects to guide dissemination strategies.
@@ 3.13.6 Operationalizing Metacognitive Verbalization and Learning-by-Teaching Processes
Metacognitive verbalization—students’ spoken articulation of cognitive processes while teaching robots—serves as both a learning mechanism and a measurable outcome. Operationalization requires structured coding schemes distinguishing utterance types: concept explanations articulating disciplinary principles, strategy statements describing problem-solving procedures, monitoring comments reflecting self-assessment of understanding, regulatory strategies for managing cognitive resources, and socio-emotional utterances [39]. Systematic time sampling (e.g., coding first and last five minutes of sessions, problem-explanation segments) yields representative estimates while managing coding burden.
Linking metacognitive behaviors to learning outcomes through mediation models tests the learning-by-teaching mechanism: increases in explanation quality or metacognitive verbalization frequency should mediate gains on curriculum-aligned assessments if the teaching-to-learn hypothesis holds [39]. This analytic approach mirrors cooperative learning research documenting that student talk quality mediates peer learning effects. Because executive function and metacognitive capacity develop across elementary grades, analyses should include EF screening or developmental covariates to contextualize verbalization patterns and test for differential responsiveness by age or grade [45].

### v2

@@ 3.12.7 Privacy, Ethics, and Equity Considerations
Sensor-rich MR systems collecting granular behavioral data from elementary students raise substantial privacy concerns requiring proactive governance [DOI_14763_2021_4_1610, DOI_33225_pec_23_81_501]. Critical scholars warn that immersive technologies risk creating surveillance infrastructures if data collection, retention, and usage policies lack transparency and community oversight. Recommendations include limiting data collection to pedagogically essential streams, anonymizing identifiable information through pseudonymization and aggregation, obtaining informed parental consent with clear communication about data purposes, and designing analytic outputs as interpretable teaching tools rather than student monitoring systems [DOI_14763_2021_4_1610, DOI_33225_pec_23_81_501, DOI_2478_amns_2023_2_01427].
@@ 3.12.8 Research Priorities and Knowledge Gaps
The field faces three critical knowledge gaps demanding targeted research investment. First, an empirical gap exists in validated XR implementations specifically designed for and evaluated with elementary STEM learners [DOI_1111_bjet_13049, DOI_1089_cyber_2023_0411]. Most rigorous effectiveness trials target adults or secondary students; evidence for elementary populations—particularly for integrated cross-disciplinary STEM learning—remains thin.
Second, a methodological gap persists in training datasets for machine learning models: publicly available multimodal behavioral datasets from children using XR systems for STEM learning do not exist, limiting researchers’ ability to develop and validate analytic algorithms or investigate generalizability across device platforms [DOI_1109_jtehm_2023_3335608, DOI_3390_buildings14020385].
Third, a pedagogical gap concerns adaptive scaffolding and teacher orchestration: research documenting effective real-time adaptation strategies based on behavioral analytics, and teacher-facing dashboard designs supporting instructional decision-making in XR-enhanced classrooms, remains urgently needed [DOI_1002_ase_2146, DOI_1111_jcal_13059].
@@ 3.12.9 Practical Recommendations for Pilot Implementations
Institutions considering XR pilots for elementary STEM should follow staged implementation guidance [DOI_1111_bjet_13049, DOI_33086_cej_v6i2_6105]. Begin with small-scale deployments (1-2 classrooms, 4-6 week instructional units) aligned closely with existing STEM curricula, co-designing activities with classroom teachers to ensure pedagogical coherence. Implement clear behavioral signal tracking from initial pilots to enable model validation and iterative refinement [DOI_3390_buildings14020385]. Prioritize interpretable analytics—visualizations teachers can understand and act upon—while maintaining transparent data governance with documented consent, storage, and usage policies [DOI_33225_pec_23_81_501, DOI_14763_2021_4_1610]. Evaluate through convergent mixed methods combining quantitative metrics with teacher interviews and classroom observations, sharing findings through practitioner networks and research publications to build the community knowledge base [DOI_1089_cyber_2023_0411, DOI_1111_bjet_13049].
Critical gap: XR/spatial-computing technologies provide compelling theoretical affordances for embodied, spatially-grounded integrated STEM learning in elementary contexts, with modern devices generating rich behavioral data streams enabling sophisticated learning analytics. However, the evidence base remains dominated by health professions training and university implementations; rigorous empirical validation specifically targeting elementary STEM classrooms—particularly for cross-disciplinary integrated learning supported by conversational AI—is lacking. Moreover, critical privacy, equity, and pedagogical integration challenges require substantial investigation before XR can move from promising prototype to validated, equitably-deployed elementary STEM learning infrastructure [DOI_3390_encyclopedia3020026, DOI_1109_tlt_2023_3324843, DOI_2478_amns_2023_2_01427, DOI_33086_cej_v6i2_6105, DOI_3390_smartcities7010001, DOI_14763_2021_4_1610].
@@ 3.13 Educational Robots and Teachable Agents: Learning-by-Teaching Paradigms for Elementary STEM
Educational robots and teachable agents represent a distinct embodied AI paradigm where elementary students learn by teaching, explaining concepts to artificial tutees rather than receiving instruction from AI tutors. This pedagogical inversion—positioning the child as expert and the AI as learner—activates metacognitive processes, explanatory elaboration, and self-monitoring that align with cooperative learning mechanisms documented to enhance elementary learning outcomes [DOI_1177_2332858420986211]. However, while cooperative peer learning shows robust meta-analytic evidence in elementary mathematics and the theoretical parallels between peer teaching and AI-tutee interaction are compelling, direct experimental evaluations of educational robots or teachable agents deployed in elementary STEM classrooms remain limited, necessitating careful synthesis of adjacent literatures to guide future empirical work [DOI_1111_ssm_12660, DOI_31756_jrsmte_531, DOI_1111_ssm_12603].
@@ 3.13.1 Theoretical Foundation: Cooperative Learning as Analogue for Tutee Interactions
Meta-analytic synthesis of elementary mathematics programs demonstrates that designs incorporating cooperative learning produce meaningful achievement gains, establishing peer explanation and collaborative problem-solving as core mechanisms linking social interaction to cognitive growth [DOI_1177_2332858420986211]. When students articulate their reasoning to peers, they externalize thought processes, identify gaps in understanding, refine explanations through iterative dialogue, and consolidate knowledge through teaching acts. These mechanisms provide direct theoretical justification for learning-by-teaching interventions with AI tutees: if explaining to human peers enhances learning, then explaining to artificial agents designed to elicit similar explanatory discourse should produce analogous cognitive benefits.
The pedagogical value of teaching others extends beyond simple rehearsal. Effective peer teaching requires monitoring the tutee’s understanding, adjusting explanations based on feedback, and scaffolding learning through strategic questioning—all metacognitive and regulatory processes that deepen the tutor’s own comprehension [DOI_1177_2332858420986211]. Educational robots and teachable agents can be designed to systematically elicit these processes through programmed responses signaling confusion, requesting clarification, or demonstrating partial understanding, thereby creating structured opportunities for elementary students to engage in explanatory and metacognitive verbalization within STEM learning contexts.
@@ 3.13.2 Implementation Realities: Teacher Preparation and Classroom Feasibility
Elementary classrooms present unique implementation constraints fundamentally shaping technology deployment feasibility. Teachers are typically subject generalists rather than STEM specialists, exhibiting variable content knowledge and pedagogical confidence across mathematics, science, engineering, and computational domains [DOI_31756_jrsmte_531, DOI_1111_ssm_12603]. Professional development supporting elementary STEM technology integration must therefore address both content knowledge and pedagogical practice, recognizing that teacher STEM identity, beliefs about technology, and confidence in orchestrating novel activities mediate implementation fidelity and sustainability.
Research on preservice and in-service elementary teacher preparation for integrated STEM instruction emphasizes co-design approaches where teachers participate in developing activities, scripting robot prompts, and establishing classroom routines before deployment [DOI_31756_jrsmte_531, DOI_1111_ssm_12603]. This participatory design philosophy treats teachers as essential partners whose practical wisdom about classroom management, developmental appropriateness, and curricular alignment ensures that robot-tutee interventions integrate smoothly into instructional sequences rather than disrupting established routines. Studies documenting effective elementary STEM teacher preparation highlight the importance of modeling exemplar implementations, providing scripted facilitation guides, and creating professional learning communities where teachers share strategies for managing technological and pedagogical challenges [DOI_31756_jrsmte_531, DOI_1111_ssm_12603].
@@ 3.13.3 Measurement Framework: Constructs, Instruments, and Developmental Considerations
Rigorous evaluation of learning-by-teaching robot interventions in elementary contexts requires multimodal assessment spanning cognitive learning outcomes, interaction processes, socio-emotional effects, and implementation fidelity [DOI_1111_ssm_12660, DOI_1177_1534508420909527]. Cognitive outcomes should employ curriculum-aligned measures with established validity evidence for the targeted grade range, following psychometric best practices for evidence-based validity and reliability reporting [DOI_1111_ssm_12660]. Recent work validating elementary mathematics assessment instruments illustrates necessary procedures: item analysis, internal consistency documentation, construct validity through correlations with established measures, and examination of measurement invariance across grade bands to ensure scores are interpretable equivalently for K-2 versus grades 3-5 students [DOI_1111_ssm_12660, DOI_1177_1534508420909527].
Interaction process measures capture the explanatory and metacognitive mechanisms hypothesized to mediate learning gains when students teach robots. Systematic sampling and coding of student-robot dialogue should prioritize: (1) concept explanations where students articulate domain principles, (2) strategy statements describing problem-solving approaches, (3) metacognitive verbalizations reflecting planning, monitoring, and evaluation, and (4) scaffolding behaviors such as questioning and prompting [DOI_1177_2332858420986211]. These codes align directly with cooperative learning literature documenting that student talk—particularly high-quality elaborations—predicts achievement gains, providing mechanistic justification for treating verbal interaction quality as a primary mediator variable [DOI_1177_2332858420986211].
Socio-emotional and behavioral outcomes require validated screening instruments appropriate for elementary populations. The Social Skills Improvement System Performance Screening Guide (SSIS PSG) demonstrates meta-analytic support as an elementary screening tool detecting behavioral and academic concerns, making it suitable for monitoring unintended effects of robot interventions on peer dynamics or classroom behavior [DOI_1177_1534508420926584]. Studies documenting structural validity and measurement invariance of socio-emotional instruments across grade clusters emphasize the importance of selecting scales whose psychometric properties support valid interpretation across the elementary range [DOI_1177_1534508420909527].
Implementation fidelity measurement captures teacher professional development participation, classroom enactment logs documenting session completion and adherence to protocols, and teacher self-reported content confidence before and after implementation [DOI_31756_jrsmte_531, DOI_1111_ssm_12603]. Because teacher knowledge and receptivity mediate technology adoption in elementary settings, measuring PD dose, teacher beliefs, and fidelity enables researchers to model these factors as moderators of student outcomes and to identify conditions supporting successful implementation [DOI_31756_jrsmte_531, DOI_1111_ssm_12603].
@@ 3.13.4 Data Collection Protocol and Analytic Strategies
A comprehensive data collection protocol for educational robot-tutee studies in elementary STEM integrates baseline and outcome cognitive testing, behavioral and socio-emotional screening, observational sampling of student-robot discourse, teacher implementation logs, and system telemetry from robot interactions [DOI_1111_ssm_12660, DOI_1177_2332858420986211]. Cognitive assessments should employ standardized or researcher-developed instruments with demonstrated validity, reporting pre-post effect sizes, growth metrics, and psychometric properties to support meta-analytic synthesis [DOI_1111_ssm_12660, DOI_1177_2332858420986211].
Discourse sampling requires audio-recording student-robot interactions and applying structured coding schemes with established inter-rater reliability. Cooperative learning evidence justifies prioritizing explanatory and metacognitive codes as intermediate outcomes theoretically linked to learning gains [DOI_1177_2332858420986211]. System logs from robots provide complementary objective data: timestamped interaction events, session durations, task sequences, and robot prompts delivered. Synchronizing telemetry with audio transcripts enables triangulation—linking coded verbal behaviors to system states and timestamped learning events to support process-outcome models [DOI_1177_2332858420986211].
Analytic approaches should employ multilevel modeling reflecting the nested structure of elementary data (students within classrooms within schools), because interventions typically deploy at classroom level and teacher practices explain substantial outcome variance [DOI_31756_jrsmte_531, DOI_1111_ssm_12603]. Hierarchical models partition variance appropriately, estimate classroom-level moderators (PD fidelity, teacher STEM identity), and support generalization to broader populations. Mediation analyses testing whether interaction processes (explanation quality, metacognitive verbalization frequency) mediate the robot intervention’s effect on learning outcomes provide mechanistic evidence aligning with cooperative learning theory [DOI_1177_2332858420986211].
Grade-band stratification and tests of measurement invariance address developmental heterogeneity across elementary years. Because working memory, executive function, and metacognitive capacity evolve substantially from kindergarten through grade 5, analyses should examine whether treatment effects, process mechanisms, or measurement properties differ across K-2 versus grades 3-5 clusters [DOI_1177_1534508420909527, DOI_1037_dev0001557]. Executive function screening or age/grade covariates help contextualize metacognitive verbalization patterns and test for differential responsiveness by developmental level.
@@ 3.13.5 Staged Research Pathway: Feasibility, Efficacy, Effectiveness
A methodologically rigorous research program for educational robots in elementary STEM should progress through staged designs balancing scientific control with ecological validity [DOI_31756_jrsmte_531, DOI_1111_ssm_12603, DOI_1177_2332858420986211]. Feasibility studies establish that robots integrate into classroom routines without excessive disruption, that teachers can implement activities with reasonable PD support, that audio/video capture and telemetry synchronization function reliably, and that instruments yield interpretable data. Teacher-led pilots allow iterative refinement of robot scripts, facilitation guides, and data collection procedures before scaling to formal trials [DOI_31756_jrsmte_531].
Efficacy trials employ cluster randomization or well-matched quasi-experimental designs assigning classrooms or schools to intervention versus comparison conditions, with comprehensive pre-post measurement, process sampling, and fidelity monitoring. Power calculations should account for classroom-level assignment and intraclass correlation typical of elementary settings [DOI_1177_2332858420986211, DOI_1111_ssm_12660]. Cooperative learning meta-analyses provide benchmarks for expected effect sizes and variance structures informing power analysis.
Effectiveness studies examine scalability and sustainability under real-world conditions, explicitly measuring teacher adoption patterns, PD scalability, and contextual moderators such as school learning models (in-person, hybrid, remote instruction) that influence technology access and supervision [DOI_1108_ils-04-2020-0136]. Pragmatic effectiveness trials prioritize external validity and generalizability, documenting implementation variation and subgroup effects to guide dissemination strategies.
@@ 3.13.6 Operationalizing Metacognitive Verbalization and Learning-by-Teaching Processes
Metacognitive verbalization—students’ spoken articulation of cognitive processes while teaching robots—serves as both a learning mechanism and a measurable outcome. Operationalization requires structured coding schemes distinguishing utterance types: concept explanations articulating disciplinary principles, strategy statements describing problem-solving procedures, monitoring comments reflecting self-assessment of understanding, regulatory strategies for managing cognitive resources, and socio-emotional utterances [DOI_1177_2332858420986211]. Systematic time sampling (e.g., coding first and last five minutes of sessions, problem-explanation segments) yields representative estimates while managing coding burden.
Linking metacognitive behaviors to learning outcomes through mediation models tests the learning-by-teaching mechanism: increases in explanation quality or metacognitive verbalization frequency should mediate gains on curriculum-aligned assessments if the teaching-to-learn hypothesis holds [DOI_1177_2332858420986211]. This analytic approach mirrors cooperative learning research documenting that student talk quality mediates peer learning effects. Because executive function and metacognitive capacity develop across elementary grades, analyses should include EF screening or developmental covariates to contextualize verbalization patterns and test for differential responsiveness by age or grade [DOI_1037_dev0001557].